You own a five-year, 6% coupon bond and the current YTM is 5%. If the YTM immediately increases to 6%, what is the percentagePlease explain throughly so I may understand.
Answer #1
Step-1:Calculation of current price
Current price =-pv(rate,nper,pmt,fv)
= $ 1,043.76
Where,
rate = 2.50%
nper = 10
pmt = $ 30.00
fv = $ 1,000.00
Step-2:Calculation of changed price
Changed Price =-pv(rate,nper,pmt,fv)
= $ 1,000.00
Where,
rate = 3.00%
nper = 10
pmt = $ 30.00
fv = $ 1,000.00
Step-3:Calculation of percentage change in the price of bond
Percentage change = (P1-P0)/P0 Where,
= (1000-1043.76)/1043.76 P0 = $ 1,043.76
= -4.19% P1 = $ 1,000.00